### v2

@@ 5.1.3 Data Preprocessing
5.2 Comparative Analysis of Noise-Robust Multi-Modal Coupling Signal Estimation Methods in Noisy Environments
@@ 5.2.1 Experimental Setup
Hospital ambient noise from [3] was added to PCG signals at SNR levels (5-30 dB), including equipment noise (ventilators, monitors) and environmental noise (conversations, movement). Performance metrics included Mean Squared Error (MSE), Pearson Correlation Coefficient (PCC), Spectral Coherence (SC), and Clinical Feature Preservation (CFP). The CFP metric [14] quantifies preservation of diagnostically relevant features.
Table 2: Comparison of estimation methods across noise levels.
@@
MSE PCC MSE PCC
30 dB Deconvolution 0.013 0.995 0.015 0.992
Tikhonov 0.010 0.996 0.012 0.994
Wiener 0.009 0.996 0.011 0.995
Sparsity 0.008 0.997 0.010 0.995
NMCSE 0.007 0.997 0.008 0.996
@@
NMCSE 0.336 0.650 0.359 0.632
5.2.2 Results and Analysis
Table 2 shows NMCSE consistently outperforms alternative methods. Spectral coherence analysis revealed NMCSE excels at preserving mid-frequency components (20-100 Hz) critical for cardiac mechanical events, with 31% higher coherence at 5 dB SNR.
CFP evaluation showed NMCSE preserved 87% of clinically relevant features at 10 dB compared to 73%, 68%, and 64% for sparsity-based, Wiener, and Tikhonov methods.
5.2.3 Spectral Coherence Analysis
To understand the frequency-domain advantages of NMCSE, we performed spectral coherence analysis comparing reference coupling signals with those estimated by different methods. Fig. 4A shows that NMCSE maintains significantly higher coherence with the reference signal in the mid-frequency range (10-100 Hz), with approximately 31% higher coherence compared to deconvolution. This frequency band contains critical mechanical cardiac events, including the first (S1) and second (S2) heart sounds.
The coherence difference map (Fig. 4B) demonstrates how NMCSE’s advantage persists across varying noise levels, with the most pronounced benefits in the mid-frequency band. As SNR decreases, NMCSE’s robustness becomes increasingly valuable, particularly at lower SNR levels (5-10 dB) typical in clinical environments.
@@ 5.2.3 Spectral Coherence Analysis
Figure 4: Spectral coherence analysis of coupling signal estimation methods. (A) Magnitude squared coherence between reference and estimated coupling signals (10dB SNR), highlighting NMCSE’s superior preservation of mid-frequency components. (B) Coherence difference map across frequencies and noise levels. (C) Average coherence preservation by frequency band (10dB SNR), quantifying NMCSE’s 28% improvement in mid-frequency preservation.
Quantitative analysis (Fig. 4C) confirms that NMCSE preserves 28% more coherence in the mid-frequency band compared to deconvolution at 10dB SNR. This explains why NMCSE-estimated coupling signals retain more diagnostically relevant features. At 10dB SNR, NMCSE preserves 87% of clinically relevant features compared to 73%, 68%, and 64% for sparsity-based, Wiener, and Tikhonov methods, respectively.
The frequency-specific performance improvement demonstrated by NMCSE aligns with our theoretical analysis in Section 3.7, confirming that our distribution-matching approach effectively mitigates noise amplification across the frequency spectrum, with particularly strong benefits in physiologically significant frequency bands.
Table 3: Impact of NMCSE parameters ($\alpha$, $\beta$) on performance metrics at SNR = 10 dB.
Metric Varying $\alpha$ (with $\beta=0.1$) Varying $\beta$ (with $\alpha=1.0$)
$\alpha=0.01$ $\alpha=0.1$ $\alpha=0.5$ $\alpha=1.0$ $\beta=0.01$ $\beta=0.1$ $\beta=0.5$ $\beta=1.0$
MSE 0.289 0.220 0.190 0.177 0.186 0.177 0.184 0.193
PCC 0.762 0.810 0.835 0.825 0.822 0.825 0.818 0.812
CFP (%) 79.3 83.5 85.6 87.2 84.9 87.2 85.7 84.5
5.2.4 Parameter Sensitivity Analysis
Table 3 shows that optimal performance is achieved at $\alpha=1.0$ and $\beta=0.1$, which is consistent with our theoretical analysis. Insufficient amplitude weighting ($\alpha<0.5$) leads to degraded performance with MSE increasing by up to 63% at $\alpha=0.01$. Similarly, performance degrades when temporal weighting ($\beta$) exceeds 0.1, with a 9% increase in MSE at $\beta=1.0$. Within the tested parameter range, NMCSE shows moderate sensitivity to parameter selection, with performance variations of less than 10% when parameters remain within 50% of their optimal values. The optimal $\alpha/\beta$ ratio remained consistent across different noise levels, suggesting that once calibrated, the same parameter configuration can be used across varying clinical conditions.
@@ 5.2.4 Parameter Sensitivity Analysis
5.3 Multi-modal CVD Detection with ECG, PCG and the NMCSE-Estimated Coupling Signal Using TSFE Network
@@
We evaluated the effectiveness of integrating the NMCSE-estimated coupling signal for CVD detection through two sets of experiments: an ablation study examining the contribution of each modality and a comparison with state-of-the-art multi-modal approaches.
5.3.1 Ablation Experiment
@@
Table 4: Evaluation metrics for different input configurations.
Input Configuration Accuracy (%) Specificity (%) Sensitivity (%) AUC
Single ECG 81.23 77.25 85.27 0.86
Single PCG 84.71 81.92 89.87 0.89
NMCSE-estimated coupling 90.25 87.48 94.13 0.93
ECG + PCG 93.85 93.02 94.67 0.96
All three modalities 97.38 97.15 97.92 0.98
As shown in Table 4, among single-modal inputs, the NMCSE-estimated coupling signal achieved the highest performance (90.25% accuracy, 0.93 AUC), outperforming both ECG (81.23%) and PCG (84.71%) alone. This demonstrates that the coupling signal effectively captures the cross-modal relationship between electrical and mechanical cardiac activities.
@@
Table 5: Performance comparison with existing methods.
@@
5.3.2 Comparison with Existing Multi-Modal Methods
We compared our approach with state-of-the-art multi-modal methods for CVD detection, summarized in Table 5.
Our approach outperforms existing methods across most metrics, achieving the highest overall accuracy (97.38%), specificity (97.15%), and AUC (0.98). Notably, our method surpasses Sun et al. [22], which also utilizes coupling signals but relies on deconvolution-based estimation. This performance difference (97.38% vs. 95.38% accuracy) directly demonstrates the advantage of using NMCSE for robust coupling signal estimation in noisy environments.
6 Conclusion
This paper introduced Noise-Robust Multi-Modal Coupling Signal Estimation, which reformulates ECG-PCG coupling signal estimation through optimal transport theory. By shifting from ill-posed deconvolution to well-posed distribution matching, NMCSE offers a theoretically sound framework that jointly optimizes amplitude and temporal alignment while mitigating noise amplification.
Experiments on PhysioNet Dataset with realistic hospital noise demonstrated NMCSE’s superior robustness, reducing estimation errors by approximately 30% compared to conventional methods including Tikhonov regularization, Wiener filtering, and sparsity-based approaches. When integrated with our Temporal-Spatial Feature Extraction network for multi-modal CVD detection, our approach achieved 97.38% accuracy and 0.98 AUC, outperforming state-of-the-art methods.
The noise robustness of NMCSE addresses a critical gap between theoretical possibility and clinical reality in coupling signal-enhanced diagnostic frameworks. By enabling reliable performance under realistic noise conditions, our method paves the way for wider adoption of multi-modal cardiac analysis in clinical settings. Future work will focus on real-time implementation and validation across diverse patient populations with varying cardiac conditions to further establish its clinical utility.

### v3

@@ V-A4 Data Preprocessing
V-B Comparative Analysis of Noise-Robust Multi-Modal Coupling Signal Estimation Methods in Noisy Environments
@@ V-B1 Experimental Setup
Hospital ambient noise from [16] was added to PCG signals at different SNR levels, including equipment noise (ventilators, monitors) and environmental noise (conversations, movement). Performance metrics included Mean Squared Error (MSE), Pearson Correlation Coefficient (PCC), Spectral Coherence (SC), and Clinical Feature Preservation (CFP). The CFP metric [27] quantifies preservation of diagnostically relevant features.
TABLE II: Performance of different estimation methods under varying SNRs and noise types. Best results are highlighted.
@@
MSE $\downarrow$ PCC $\uparrow$ MSE $\downarrow$ PCC $\uparrow$
30 dB Deconvolution 0.013 0.921 0.015 0.912
Tikhonov 0.010 0.925 0.012 0.924
Wiener 0.009 0.945 0.011 0.943
Sparsity 0.008 0.962 0.010 0.955
NMCSE 0.0065 0.997 0.0075 0.996
@@
NMCSE 0.310 0.670 0.325 0.655
V-B2 Results and Analysis
Table II demonstrates that NMCSE consistently outperforms all baselines across different SNR levels and noise types. It achieves the lowest mean squared error (MSE) and highest Pearson correlation coefficient (PCC) in every setting. At 5 dB SNR, NMCSE reduces MSE by up to 24% and improves PCC by 7.3% compared to the best-performing baseline. These improvements are more pronounced under severe noise, indicating the superior robustness and reliability of NMCSE in challenging environments.
V-B3 Spectral Coherence Analysis
We conducted spectral coherence analysis to evaluate how well each method preserves frequency-specific structure of the coupling signal. As shown in Fig. 4A, NMCSE achieves substantially higher coherence with the reference signal in the mid-frequency range (10–100 Hz), improving coherence by 31% over deconvolution. This range is clinically important as it captures key mechanical events like S1 and S2 heart sounds.
@@ V-B3 Spectral Coherence Analysis
Figure 4: Spectral coherence analysis of NMCSE method. (A) Magnitude-squared coherence between reference and estimated signals at 10 dB SNR, showing NMCSE’s superior preservation in the mid-frequency band (10–100 Hz). (B) Coherence difference map across frequencies and SNRs, highlighting advantages of NMCSE under noise. (C) Average coherence preservation by frequency band at 10 dB SNR, with NMCSE achieving a 28% gain in the mid-frequency band.
The coherence difference map in Fig. 4B confirms that NMCSE consistently outperforms across noise levels, with the strongest gains in the mid-frequency band under lower SNRs—conditions typical of real-world clinical settings.
Fig. 4C further quantifies this advantage, showing a 28% increase in average coherence preservation in the mid band at 10 dB SNR. These results support the conclusion that NMCSE not only improves noise robustness but also more effectively captures physiologically relevant dynamics.
TABLE III: Effect of NMCSE hyperparameters $\alpha$ and $\beta$ on performance at 10 dB SNR. Bold indicates best results.
Metric Varying $\alpha$ (fixed $\beta=0.1$) Varying $\beta$ (fixed $\alpha=1.0$)
0.01 0.1 0.5 1.0 0.01 0.1 0.5 1.0
MSE $\downarrow$ 0.289 0.220 0.190 0.177 0.186 0.177 0.184 0.193
PCC $\uparrow$ 0.762 0.810 0.835 0.825 0.822 0.825 0.818 0.812
CFP (%) $\uparrow$ 79.3 83.5 85.6 87.2 84.9 87.2 85.7 84.5
V-B4 Parameter Sensitivity Analysis
Table III indicates that optimal performance is achieved at $\alpha=1.0$, $\beta=0.1$, aligning with our theoretical expectations. Low amplitude weighting ($\alpha<0.5$) significantly degrades performance, with MSE increasing by up to 63% at $\alpha=0.01$. Similarly, overly strong temporal weighting ($\beta>0.1$) results in up to a 9% MSE increase. Within 50% of the optimal values, performance remains stable (variation $10\%$), and the optimal $\alpha/\beta$ ratio generalizes well across noise levels—highlighting the robustness of the parameter setting in clinical environments.
V-C Intra-State Consistency Evaluation
Previous experiments validated the robustness of coupling signal estimation under externally added ambient noise, such as hospital equipment sounds and environmental interference. While these scenarios simulate common clinical disturbances, they do not fully capture the internal physiological noise introduced by the subject’s own motion—such as body sway or respiration—that naturally occurs during physical activities.
To further investigate robustness under such real-world conditions, we design a complementary experiment based on intra-state consistency, using the EPHNOGRAM dataset [17]. This experiment specifically targets the high-intensity activity condition (e.g., stair climbing), where physiological noise and movement artifacts are most prominent.
For each subject, we first estimate a reference coupling signal $h_{\text{ref}}$ from a clean resting-state ECG-PCG segment. Since resting signals are minimally affected by motion, $h_{\text{ref}}$ serves as a reliable subject-specific baseline that reflects the core electromechanical transformation. We then select two non-overlapping 10-second segments from the subject’s high-intensity activity recording and estimate two coupling signals, $h^{(A)}$ and $h^{(B)}$, using each method independently.
We evaluate consistency in two ways:
  • Reference Consistency: We compare $h^{(A)}$ and $h^{(B)}$ to the rest-state reference $h_{\text{ref}}$ using MSE and PCC. A robust method should produce activity-state estimates that remain structurally close to the baseline signal despite motion-induced noise.
  • Mutual Consistency: We also compute MSE and PCC between $h^{(A)}$ and $h^{(B)}$ directly, to assess internal consistency of the method’s output under repeated estimation within the same noisy state.
TABLE IV: Intra-State Consistency Evaluation under Physiological Noise based on EPHNOGRAM dataset
Method Ref. Consistency Mutual Consistency
MSE $\downarrow$ PCC $\uparrow$ MSE $\downarrow$ PCC $\uparrow$
Deconvolution 0.312 0.622 0.128 0.682
Tikhonov 0.284 0.654 0.115 0.715
Wiener 0.261 0.673 0.104 0.753
Sparsity 0.245 0.688 0.092 0.770
NMCSE (Ours) 0.186 0.734 0.063 0.842
This design enables the evaluation of algorithmic stability under real physiological noise. High intra-state consistency indicates the method can suppress internal perturbations while preserving consistent electromechanical patterns. Such stability is crucial for wearable or ambulatory applications, where noise is unpredictable and labels are unavailable.
Result Analysis: As shown in Table IV, NMCSE achieves the best performance across all metrics, with the lowest error and highest similarity to the rest-state reference, and strong consistency between repeated estimates. This suggests that NMCSE is robust to physiological noise and produces stable outputs within the same activity state. In contrast, deconvolution-based methods exhibit greater variability, reflecting their sensitivity to transient signal fluctuations.
@@ V-C Intra-State Consistency Evaluation
V-D Multi-modal CVD Detection with ECG, PCG and the NMCSE-Estimated Coupling Signal Using TSFE Network
@@
We evaluated the effectiveness of our method for CVD detection through two sets of experiments: an ablation study examining the contribution of each modality and a comparison with state-of-the-art multi-modal approaches.
V-D1 Ablation Experiment
@@
TABLE V: CVD detection performance with different input configurations.
Input Configuration Accuracy (%) Specificity (%) Sensitivity (%)
Single ECG 81.23 77.25 85.27
Single PCG 84.71 81.92 89.87
NMCSE-estimated Coupling 90.25 87.48 94.13
ECG + PCG 93.85 93.02 94.67
All Three Modalities 97.38 97.15 97.92
As shown in Table V, among single-modal inputs, the NMCSE-estimated coupling signal achieved the highest performance, outperforming both ECG and PCG alone. This demonstrates that the coupling signal effectively captures the cross-modal relationship between electrical and mechanical cardiac activities.
@@
TABLE VI: Performance comparison with existing methods.
@@
V-D2 Comparison with Existing Multi-Modal Methods
We compared our method with several state-of-the-art multi-modal approaches for CVD detection, as summarized in Table VI.
Our method achieves the highest overall performance. Notably, it outperforms Sun et al. [11], which also incorporates coupling signals but relies on deconvolution-based estimation, resulting in reduced robustness under real-world noise.
Recent Transformer-based methods such as CAD-ViT [22] and PACFNet [23] enhance multi-modal fusion by modeling cross-modal dependencies through stacked self-attention modules. However, these approaches do not explicitly capture the underlying electromechanical coupling between ECG and PCG and are often sensitive to noise due to their reliance on fully supervised training. In contrast, our method introduces a physiologically informed coupling estimation stage based on optimal transport, followed by a unified TSFE network. This design enables interpretable and noise-robust integration of multi-modal cardiac signals, making it more suitable for real-world clinical environments.
VI Conclusion
We propose NMCSE, a noise-robust coupling signal estimation method that formulates ECG-PCG alignment as a distribution-matching problem via optimal transport. This approach avoids the instability of deconvolution and enables joint alignment of amplitude and timing under noise. Experiments on both simulated and real-world noisy datasets show that NMCSE outperforms existing methods in estimation accuracy and stability. When integrated with a TSFE-based multi-modal network, it achieves state-of-the-art performance in CVD detection. Source code and pretrained models will be released upon acceptance.